from 2 stock solutions 18% and 42% you are to prepare 1.5 liters of a 34% solution. How many milligrams of each stock do you need?
Answer by stanbon(75887)

from 2 stock solutions 18% and 42% you are to prepare 1.5 liters of a 34% solution. How many milligrams of each stock do you need?
-----
Equation:
active + active = active
0.18x + 0.42(1.5-x) = 0.34*1.5
----
Multiply thru by 100 to get:
18x + 42*1.5-42x = 34*1.5
-24x = -18
x = 3/4 liters (amt of 18% solution needed)
---
1.5-x = 3/4 liters (amt of 42% solution needed)
==================
Cheers,
Stan H.
=============